OpenStudy (anonymous):

Show all work. Find the following product (y-9)^2

OpenStudy (jlastino):

(y-9)^2 (square the first term) + 2(product of first and second) + (square the last term)

OpenStudy (experimentx):

\((y-9)^2 = (y - 9)(y - 9)\) \(=y^2 - 9y - 9y + 81\) \(=y^2 - 18y +81\)

OpenStudy (anonymous):

thanks!